# Discord-type quantum correlations in axially symmetric spin-(1/2, $S$) systems

### v1

@@ Abstract
A mixed spin-$(1/2,S)$ system with arbitrary spin $S$ and interactions satisfying the U(1) axial symmetry is considered. Compact explicit forms of the local quantum uncertainty (LQU) and local quantum Fisher information (LQFI) are obtained directly through the entries and eigenvalues of the density matrix. Using these forms, we perform a comparative study of discord-type quantum correlations LQU and LQFI for the system at thermal equilibrium. An increase in quantum correlations with increasing spin length $S$ is discovered. Moreover, we find that as the system cools, quantum correlations can undergo one or more abrupt transitions while the temperature changes smoothly.
@@ I Introduction
Quantum information technologies are rapidly entering all aspects of human life. The current state of quantum computing is called the Noisy Intermediate-Scale Quantum (NISQ) era [1, 2, 3]. Today’s quantum computers are limited to hundreds of qubits and have relatively short coherence times and significant errors [2, 4]. In general, they are far from universal, fault-tolerant quantum computers, which are believed to make it possible to effectively simulate any processes occurring in Nature.
Nevertheless, existing quantum computers can already solve important real-world problems. In search of the quantum advantage of these devices, NISQ algorithms have been proposed for applications in various disciplines, including physics, quantum chemistry, materials science, machine learning, etc. The goal of such algorithms is to exploit the limited available resources to perform classically complex tasks.
@@
At the heart of quantum computing lie extraordinary correlations between subsystems of a quantum-mechanical system. Until the 21st century, the quantum entanglement was considered to be the only ingredient of quantum properties leading to such correlations. It manifests itself in the Einstein-Podolsky-Rosen gedanken (thought) experiment, Bell’s inequality test, quantum cryptography, superdense coding, teleportation, and so on [5, 6] (see also review articles [7, 8, 9, 10]).
Quantum entanglement was quantified in 1996, first for pure states [11, 12], and then for mixed states [13]. According to the accepted definition, the entanglement of a bipartite pure state is the von Neumann entropy of either of the two subsystems. The entanglement (of formation) of a bipartite mixed state is defined as the minimum entanglement of an ensemble over all ensembles realizing the mixed state; using this definition, Wootters derived an exact formula for general two-qubit systems [14]. Another important measure of the quantum entanglement was proposed in 2002 by Vidal and Werner [15]. This measure, called negativity, originated from the Peres-Horodecki PPT-criterion [16, 17]. Moreover, many measures have been introduced so far to quantify the entanglement of bipartite and multipartite systems, each of which is efficient for specific quantum states [18, 19].
At one time it was believed that quantum entanglement was the only resource of quantum speedup in quantum computation. However, in 1998, Knill and Laflamme showed, using the model of deterministic quantum computation with one pure qubit (DQC1) [20], that the computation can achieve an exponential improvement in efficiency over classical computers even without containing much entanglement. In 2000-2001, $\rm\dot{Z}$urek et al. developed the concept of quantum discord – “a measure of the quantumness of correlations” [21, 22]. Simultaneously and independently, Vedral et al. [23, 24] proposed a measure for the purely classical correlation, which, after subtracting it from the total correlation, led to the same amount of quantum correlation as the discord. Then, Datta et al. [25, 26] calculated discord in the Knill-Laflamme DQC1 model and proved that it scales with the quantum efficiency, while entanglement remains vanishingly small throughout the computation. This attracted a lot of attention to the new measure of quantum correlation [27, 28, 29, 30, 31, 32].
Quantum discord and entanglement are the same for the pure quantum states. However, discord can exist in separable mixed states, i.e., when quantum entanglement is identically equal to zero. The set of separable states possesses a nonzero volume in the whole Hilbert space of a system [33] (it is a necessary condition for the arising of entanglement sudden death (ESD) effect [34]), whereas the set of states with zero discord, vise versa, is negligibly small [35]. This circumstance alone sharply distinguishes discord from entanglement. Moreover, numerous theoretical and experimental investigations of different quantum systems have clearly shown that while the quantum entanglement and discord measure the same thing – the quantum correlation – but as a matter of fact, discrepancies in quantitative and even qualitative behavior are large [36, 37, 38]. Discord and entanglement behave differently even for the simplest mixed states, namely the Werner and Bell-diagonal ones (see, e.g., [39]). This has led many to talk about entanglement and discord as different types of quantum correlations.
The subsequent proposals with more and more new measures of quantum correlations [40, 41] caused a dilemma: should each measure be attributed to its own correlation, or should it be argued that there is only one quantum correlation, but the methods of describing it are not adequate? The physicists community now prefers to talk about entanglement and discord-like quantum correlations [42]. As is customary for brevity, we will also refer to the various measures of quantum correlation as “quantum correlations”. Nevertheless, the quantum correlation is one, but now there are only different measures for it, which are still imperfect. We would like to have such measures of quantum correlations that would be useful for estimating the values of the speedup of quantum computing, the efficiency of quantum heat engines, etc. Let the measures be different, but the results must be the same. As, for example, there are various formulations of quantum mechanics which differ dramatically in mathematical and conceptual overview, yet each one makes identical predictions for all experimental results [43]. The existing difficulties only increase interest in quantum correlations because “who knows, studying the nature of correlations in the world around us might even help us catch a glimpse of the theory that comes to supersede quantum physics” [29].
@@
The concept of quantum discord is based on local measurements and optimization that were applied to the mutual information. This idea was then used for other physical or information quantities, and to date, a large number of discord-like measures of quantum correlation have been introduced, such as quantum deficit, geometric quantum discord, measurement induced disturbance and so on [29, 40, 41]. The local quantum uncertainty (LQU) and local quantum Fisher information (LQFI), based respectively on the Wigner-Yanase and quantum Fisher information, belong to this type of measures [44, 45].
@@
Nonclassical correlations in various special cases of the mixed spin-$(1/2,S)$ model were considered in a number of works. For example, the quantum entanglement, quantified by negativity, was investigated in Refs. [51, 52, 53, 54, 55]. Recently, two studies have been carried out on quantum correlations in hybrid qubit-qutrit systems using negativity, as well as LQU and LQFI [56, 57]. In the paper [58], we provided exact formulas of LQU and LQFI for general qubit-qubit X states. Extending this approach to arbitrary spin $S$, we evaluate here the exact analytical expressions for LQU and LQFI in the case of $(1/2,S)$ axisymmetric states and examine the behavior of thermal quantum correlations for different choices of Hamiltonian parameters.
@@
The present paper is structured as follows. In the next section, we establish the general form of the axially symmetric Hamiltonian and density matrix. Section III is devoted to the diagonalization of the density matrix. In Section IV, we derive formulas suitable for actually calculating the LQU and LQFI correlations. The spin-(1/2,$S$) system at thermal equilibrium is considered in Sect. V. Section VI includes the numerical results and discussion. Finally, concluding remarks close this paper.
II Hamiltonian and density matrix
[…]
@@ II Hamiltonian and density matrix
The most general matrix which commutes with the $z$-component of total spin (1) can be written as
${\cal A}=\left(\begin{array}[]{cccccccccc}a_{00}&&&&&&&&&\\ &a_{11}&0&\ldots&0&a_{1,2S+1}&0&\ldots&0&\\ &0&a_{22}&\ldots&0&0&a_{2,2S+2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&a_{2S,2S}&0&0&\ldots&a_{2S,4S}&\\ &a_{2S+1,1}&0&\ldots&0&a_{2S+1,2S+1}&0&\ldots&0&\\ &0&a_{2S+2,2}&\ldots&0&0&a_{2S+2,2S+2}&\ldots&0&\\ &\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots&\\ &0&0&\ldots&a_{4S,2S}&0&0&\ldots&a_{4S,4S}&\\ &&&&&&&&&a_{4S+1,4S+1}\end{array}\right).$ (2)
This matrix has a block-diagonal structure $(1\times 1)\oplus(4S\times 4S)\oplus(1\times 1)$, where the interior $4S\times 4S$ subblock is sparse, and its nonzero entries are located only on the main diagonal and two sub-diagonals. Remarkably, the set of matrices of such a type are algebraically closed: their sums and products preserve the same form. Below, for brevity, such matrices generated by the axial symmetry will be called $\cal A$ matrices.
If the matrix (2) is Hermitian, as in the case of Hamiltonians or density matrices, it may have $2(4S+1)$ real independent parameters. Expansion of the matrix ${\cal H}$ into the spin-1/2 and spin-$S$ matrices and their products and powers allows us to clarify the structure and physical meaning of interactions in the system (see below).
For a given Hamiltonian $\cal H$ in the $\cal A$ form, the density matrix $\rho$ can be obtained from solving some master equations. Let the initial state has the $\cal A$ structure and the dynamics of the system is governed, for example, by the quantum Liouville equation. Then, the $\cal A$ form of the density matrix will be preserved throughout the entire evolution process due to the above-mentioned property of the algebraic closedness of $\cal A$ matrices. The $\cal A$-form of the density matrix will also be realized for systems in thermal equilibrium, since $\rho$ and $\cal H$ are related by functional dependence. In these and similar cases, the density matrix under examination can be written as
@@
Below, we restrict ourselves to these $\cal A$ quantum states.
III Diagonal representation of the density matrix
III.1 Diagonalization of the matrix $\rho$
Let us turn to the diagonalization of matrix (3). Since the density matrix $\rho$ commutes with the diagonal matrix ${\cal S}_{z}$, it has nonzero matrix elements only between states with the same eigenvalue of ${\cal S}_{z}$. Here, ${\cal S}_{z}$ plays a role of “constant or integral of motion”; in this connection see, e.g., Chapter 1 in two books [59, 60]. As mentioned above, ${\cal S}_{z}$ has two non-degenerate and $4S$ doubly degenerate eigenvalues. Therefore, the density matrix (3) actually consists of two one-by-one and $2S$ two-by-two subblocks.111 This will be obvious if we reorder the vectors of the original basis so that the $k$th and $(2S+k)$th vectors become neighbors, then the density matrix (3) takes an explicit quasi-diagonal form: $\hat{\rho}=[p_{0},\left(\begin{array}[]{cc}a_{1}&u_{1}\\ u_{1}^{*}&a_{2S+1}\\ \end{array}\right),\ldots,\left(\begin{array}[]{cc}a_{2S}&u_{2S}\\ u_{2S}^{*}&a_{4S}\\ \end{array}\right),p_{4S+1}].$ This form is implemented for the system $(S,1/2)$. The first and $(4S+1)$th “subblocks” have sizes $1\times 1$, i.e. they are ready eigenvalues. The remaining $2S$ subblocks are $2\times 2$-dimensional, and their diagonalization is easily accomplished by directly calculating the eigenvectors of the corresponding subblocks (see, for example, [61]). Eigenvectors are constructed using the $k$th and $(2S+k)$th pair of vectors of the original basis. Indeed, a suitable fragment extracted from the inner $4S\times 4S$ block of the density matrix (3) reads
and then the corresponding eigenvectors are given as
where
where ${\tilde{q}}_{k}$ and ${\tilde{u}}_{k}$ ($k=1,\ldots,2S$) are defined by Eq. (6). Note that $R$ is the $\cal A$ matrix. Moreover, it is clear that $R^{\dagger}=R$ and $R^{\dagger}R=I$, so this diagonalizing transformation is both Hermitian and unitary.
Using Eqs. (3) and (7) we find
where
$p_{k,2S+k}=\frac{1}{2}\Big{(}a_{k}+a_{2S+k}\pm\sqrt{(a_{k}-a_{2S+k})^{2}+4|u_{k}|^{2}}\Big{)}$ (9)
are the eigenvalues of the density matrix (3). From here and Eq. (6) it follows that
$q_{k}=p_{k}-a_{2S+k}$ (10)
and
where as before $k=1,\ldots,2S$.
III.2 Spin matrices $\sigma_{\mu}\otimes I_{2S+1}$ in $\rho^{\prime}$-representation
Let us transform the local Pauli matrices $\sigma_{\mu}\otimes I_{2S+1}$ ($\mu=x,y,z$) of the particle $A$ into a diagonal representation of the density operator. Using Eq. (7) we get
and
Unlike the first two matrices, the latter has the $\cal A$ form. The entries $\langle m|\sigma_{\mu}\otimes I_{2S+1}|n\rangle$ of the matrices (12) – (14) will be needed in the next section.
IV Quantum correlations
IV.1 LQU
@@
Importantly, the authors [44] were able to perform optimization for qubit-qudit systems and presented the measure (18) in the simple form
${\cal U}=1-\lambda_{max}^{(W)},$ (19)
where $\lambda_{max}^{(W)}$ denotes the maximum eigenvalue of the $3\times 3$ symmetric matrix $W$ whose entries are
@@
with $\mu,\nu=x,y,z$ and $\sigma_{x,y,z}$ are the Pauli matrices as before. This opens up the possibility of real calculating LQU.
Using Eqs. (12)–(14), we find that the matrix $W$ is diagonal, $W_{yy}=W_{xx}$ and
@@
Substituting expressions (11) into Eqs. (IV.1) and (22), we arrive at the final formulas for the branches
@@
It is remarkable that both quantities ${\cal U}_{0}$ and ${\cal U}_{1}$ are expressed directly only through the matrix elements of the density matrix (3) and its eigenvalues, therefore, the same is true for the measure ${\cal U}$.
IV.2 LQFI
Quantum Fisher information (QFI) is the cornerstone of quantum estimation theory [65, 66, 67, 68, 69]. Suppose there is unitary evolution of the quantum state $\varrho=e^{iH\epsilon}\rho e^{-iH\epsilon}$ with some observable $H$, and the parameter $\epsilon$ wanted to be estimated. The QFI can be defined as follows
@@ IV.2 LQFI
${\cal F}=1-\lambda_{max}^{(M)},$ (30)
where $\lambda_{max}^{(M)}$ is the largest eigenvalue of the real symmetric $3\times 3$ matrix $M$ with entries [72]
@@
Again using Eqs. (12)–(14), we find that the matrix $M$ is diagonal, $M_{yy}=M_{xx}$, and
@@
It is interesting to draw our attention to the analogy in the structures, on the one hand, between Eqs. (IV.1) and (32) and, on the other hand, between Eqs. (22) and (33).
@@
${\cal F}_{1}=1-4\Bigg{\{}\frac{p_{0}(a_{2S+1}p_{0}+p_{1}p_{2S+1})}{(p_{0}+p_{1})(p_{0}+p_{2S+1})}+\frac{p_{4S+1}(a_{2S}p_{4S+1}+p_{2S}p_{4S})}{(p_{2S}+p_{4S+1})(p_{4S}+p_{4S+1})}+\frac{1}{2}\sum_{k=1}^{2S-1}\Bigg{[}\frac{(a_{k}p_{k+1}+p_{k}p_{2S+k})p_{k+1}}{(p_{k}+p_{k+1})(p_{k+1}+p_{2S+k})}$
@@
$+\frac{(a_{k}p_{2S+1+k}+p_{k}p_{2S+k})p_{2S+1+k}}{(p_{k}+p_{2S+1+k})(p_{2S+k}+p_{2S+1+k})}\Bigg{]}\Bigg{\}}.$ (36)
The resulting formulas (24), (25), (35) and (IV.2) for the branches of quantum correlations are surprisingly simple, compact and very comfortable to perform quick calculations on a computer.
V System at thermal equilibrium
Let’s now turn to the consideration of a spin system in the state of thermal equilibrium with a reservoir at the temperature $T$.
V.1 Spin Hamiltonian and Gibbs density matrix
We take the Hamiltonian in the Hermitian $\cal A$ form

### v2

# Quantum correlations in general qubit-qudit axially symmetric states

@@ Abstract
Abstract. The development and improvement of analytical methods for evaluating nonclassical correlations is one of the most important tasks in quantum information science. In this paper, we investigate a mixed spin-$(1/2,S)$ system with an arbitrary spin $S$, where the interactions satisfy the U(1) axial symmetry. Analytical formulas for the local quantum uncertainty (LQU) and local quantum Fisher information (LQFI) are derived directly from the elements and eigenvalues of the density matrix. These results are then used to conduct a comparative analysis of the discord-like quantum correlations, LQU and LQFI, in the system at thermal equilibrium. The high-temperature asymptotics of both quantum correlations are found explicitly. Despite the destructive role of temperature in general, the calculations show that the quantum correlations can increase with temperature in local intervals. Under certain conditions, temperature even generates quantum correlations from uncorrelated ground states. Further, as the system cools, quantum correlations can undergo a series of abrupt transitions with a smooth temperature change. These phenomena are demonstrated for different choices of coupling parameters and spin lengths $S$.
@@ I Introduction
Quantum information technologies are rapidly entering all aspects of human life. The current state of quantum computing is called the Noisy Intermediate-Scale Quantum (NISQ) era [1, 2, 3]. Today’s quantum computers are limited to hundreds of qubits and have relatively short coherence times and significant errors [2, 4]. In general, they are far from universal, fault-tolerant quantum computers, which are believed to make it possible to effectively simulate any processes occurring in Nature. Nevertheless, existing quantum computers can already solve important real-world problems.
@@
It is believed that the speedup of quantum computers is due to quantum correlations, although there is no rigorous evidence. Until the 21st century, the quantum entanglement was considered to be the only ingredient of quantum properties leading to such correlations. It emerges in various phenomena and protocols, including the Einstein-Podolsky-Rosen thought experiment, Bell’s inequality tests, superdense coding, teleportation, quantum cryptography, and so on [5, 6].
@@
Many measures of quantum correlations have been proposed to date, with their properties now being extensively investigated through both theoretical and experimental approaches [7, 8, 9, 10, 11, 12, 13]. Quantum correlations are characterized as the difference between total and classical correlations [14, 15]. Besides entanglement, the most important among them are quantum discord, one-way quantum work deficit and some others [16, 17]. The idea of quantum discord is based on local measurements and optimization that were applied to the mutual information. This concept was subsequently extended to various physical and informational quantities, leading to the introduction of some discord-type measures of quantum correlations. Notable examples include geometric quantum discord, measurement-induced disturbance, and so forth [9, 16, 17]. The local quantum uncertainty (LQU) and local quantum Fisher information (LQFI), based respectively on the Wigner-Yanase and quantum Fisher information, belong to this type of measures [18, 19, 20].
For pure quantum states, quantum discord and entanglement coincide. However, unlike entanglement, discord can persist in separable mixed states, where entanglement is strictly zero. Notably, the set of separable states occupies a finite volume within the system’s Hilbert space [21], whereas the set of zero-discord states is vanishingly small [22]. This fundamental distinction highlights a key difference between discord and entanglement. Furthermore, Datta et al. [23, 24] evaluated discord in the Knill-Laflamme DQC1 model [25], demonstrating that it scales with quantum efficiency, even as entanglement remains negligible throughout the computation.
@@
The two-site spin-$(1/2,S)$ models have been considered in a number of works. In particular, SU(2)-invariant states were studied in Refs. [31, 32]. These states are invariant under uniform rotation of both spins, which means that the density matrix commutes with all components of the total spin. For it, many important nonclassical correlations, such as quantum entanglement in the form of negativity [31, 32], quantum entanglement of formation [33], relative entropy of entanglement [34], quantum discord [35], one-way deficit [36] and local quantum uncertainty [37] have been found. Moreover, the Holevo quantity has been also obtained in [38]. Unfortunately, SU(2)-invariant states are single-parameter and in fact correspond to only one model–a fully isotropic Heisenberg dimer.
In this paper, we consider $(1/2,S)$-systems with lower symmetry, namely with the axial symmetry, that is, when the density matrix commutes with only one component of the total spin. Unlike the SU(2) case, the axially symmetric state contains $18S+$ 1 real independent parameters (see Sec. II) and covers many physically important systems such as Heisenberg XXZ model in an inhomogeneous external magnetic field with Dzyaloshinskii-Moriya (DM) interaction, various cases of single- and two-ion anisotropies, etc. For such systems, the quantum entanglement, quantified by negativity, was investigated in Refs. [39, 40, 41].
Many closely related works in the literature have explored quantum correlations and the influence of interaction terms on quantum systems. For instance, Erol et al. [42] analyzed various entanglement measures and the local operations and classical communication (LOCC)-maximized quantum Fisher information in two-qubit systems, providing insights into quantum estimation theory. Ozaydin and Altintas [43] demonstrated how DM interaction can enhance quantum metrology by surpassing the shot-noise limit, highlighting its significance for precision measurements. Slaoui et al. [44] conducted a comparative study between quantum correlations quantifiers LQFI and LQU in the Heisenberg isotropic (and anisotropic) XY model subjected to an external magnetic field, showing that LQU is majorized by LQFI in any metrological task of the phase estimation. Haseli [45] investigated LQFI and LQU in a two-qubit Heisenberg XYZ chain with DM interaction, shedding light on their impact on quantum uncertainty quantifiers. Lian and Liu [46] extended the study of quantum Fisher information to a qubit-qutrit system in the context of Garfinkle–Horowitz–Strominger dilaton space-time, exploring relativistic effects on quantum estimation. Fedorova and Yurischev [47] examined quantum discord, LQU and LQFI in a two-spin-1/2 Heisenberg chain with both DM and Kaplan–Shekhtman–Entin-Wohlman–Aharony (KSEA) interactions, emphasizing the temperature effects and system parameters. Additionally, Mahdavifar et al. [48] studied the resilience of quantum spin fluctuations under DM interaction, revealing its role in stabilizing quantum correlations in spin systems. Also, a study of quantum correlations in hybrid $(1/2,1)$-systems was carried out using negativity, as well as LQU and LQFI [49].
Compared to our recent papers [50, 51, 52], which focused on deriving compact closed-form expressions for LQU and LQFI in general qubit-qubit X states and qubit-qutrit axially symmetric states, our present work extends the analysis to a more generalized mixed spin-$(1/2,S)$ system with arbitrary spin $S$. While the previous study [52] specifically examined a spin-$(1/2,1)$ system with ten independent interaction parameters, the current work formulates analytical expressions for LQU and LQFI in a broader class of systems satisfying U(1) axial symmetry, offering a more comprehensive framework for evaluating nonclassical correlations in a hybrid spin-$(1/2,S)$ system.
The significance and novelty of our present paper lie in its generalized formulation, which allows for the study of arbitrary spin-$(1/2,S)$ systems, thereby broadening the applicability of LQU and LQFI as tools for characterizing quantum correlations. The analytical expressions derived for these measures enable a direct and efficient evaluation of quantum correlations from the density matrix entries and eigenvalues, simplifying their computation in a wide range of physical systems. The discovery of temperature-induced quantum correlations and multiple abrupt transitions highlights new aspects of thermal quantum effects that could be relevant for quantum thermodynamics and quantum information processing. By systematically analyzing how quantum correlations behave across different temperature regimes and spin configurations, this study provides deeper physical insights into the interplay between thermal fluctuations and quantum coherence, offering potential applications in quantum materials and quantum technologies.
@@
The present paper is structured as follows. In the next section, the general form of the axially symmetric density matrix is established. In Sec. III, we derive formulas that are suitable for actually calculating LQU and LQFI correlations. The spin-(1/2,$S$) system at thermal equilibrium is considered in Sec. IV. Sec. V includes high-temperature expansions of quantum correlations, numerical results and discussion. Finally, concluding remarks in Sec. VI close the paper.
II A form of axially symmetric states
Take a system composed of two particles with spins 1/2 and $S$. Let its quantum state be invariant under transformations of the axial symmetry group U(1), consisting of rotations $R_{z}(\phi)=\exp(-i\phi{\cal S}_{z})$ around the $z$-axis on angle $\phi\in[0,\pi)$. This means that the density matrix commutes with the $z$-component of total spin, given by
@@ II A form of axially symmetric states
The most general Hermitian matrix which commutes with the $z$-component of total spin (1) can be written as
@@
This matrix has a block-diagonal structure $(1\times 1)\oplus(4S\times 4S)\oplus(1\times 1)$, where the interior $4S\times 4S$ subblock is sparse, and its nonzero entries are located only on the main diagonal and two sub-diagonals. Positive definiteness is assumed for any density matrices ($\rho\geq 0$), as well as normalization of the trace to unity (${\rm Tr}\rho=1$). In the case under consideration, this means that $p_{0},a_{1},\ldots,a_{4S},p_{4S+1}\geq 0$, $a_{k}a_{2S+k}\geq|u_{k}|^{2}$ and $p_{0}+p_{4S+1}+\sum_{i=1}^{4S}a_{i}=1$. Below, for brevity, the density matrices like (2) will be called as the axially symmetric states. Obviously, the density matrix (2) contains $18S+$ 1 free real parameters.
The transformation $R$ diagonalizing $\rho$ is presented in Appendix A. It allows us to extract the eigenvalues of the interior subblock $4S\times 4S$ of the density matrix (2), which are given as
$p_{k,2S+k}=\frac{1}{2}\Big{(}a_{k}+a_{2S+k}\pm\sqrt{(a_{k}-a_{2S+k})^{2}+4|u_{k}|^{2}}\Big{)}.$ (3)
Now, we proceed to the derivation of formulas for the quantum correlations LQU and LQFI.
III Quantum correlations for axially symmetric states
III.1 LQU
@@
Importantly, the authors [18] were able to perform optimization for qubit-qudit systems and presented the measure (7) in the following form
${\cal U}=1-\lambda_{\max}^{(W)},$ (8)
where $\lambda_{\max}^{(W)}$ denotes the maximum eigenvalue of the $3\times 3$ symmetric matrix $W$ whose entries are
@@
with $\mu,\nu=x,y,z$ and $\sigma_{x,y,z}$ are the Pauli matrices. In principle, this opens up the possibility of analytical calculating LQU by solving secular (algebraic) equation of the third degree for the matrix $W$.
However, passing under the trace symbol in Eq. (9) to the diagonal representation of the density matrix (2) and using Eqs. (49)–(51) from Appendix B, we find that the matrix $W$ for axially symmetric states is diagonal, $W_{yy}=W_{xx}$ and
@@
Substituting expressions (48) into Eqs. (III.1) and (11), we arrive at the final formulas for the branches
@@
III.2 LQFI
Quantum Fisher information (QFI) is the cornerstone of quantum estimation theory [56, 57, 58, 59]. Suppose there is a unitary evolution of the quantum state $\varrho=e^{iH\epsilon}\rho e^{-iH\epsilon}$ with some observable $H$, and the parameter $\epsilon$ needs to be estimated. Thus, the QFI can be defined as follows
@@ III.2 LQFI
${\cal F}=1-\lambda_{\max}^{(M)},$ (19)
where $\lambda_{\max}^{(M)}$ is the largest eigenvalue of the real symmetric $3\times 3$ matrix $M$ with entries [61]
@@
Again using Eqs. (49)–(51), we find that the matrix $M$ for axially symmetric states is also diagonal, $M_{yy}=M_{xx}$, and
@@
It is interesting to draw our attention to the analogy in the structures, on the one hand, between Eqs. (III.1) and (21) and, on the other hand, between Eqs. (11) and (22).
@@
${\cal F}_{1}=1-4\Bigg{\{}\frac{p_{0}(a_{2S+1}p_{0}+p_{1}p_{2S+1})}{(p_{0}+p_{1})(p_{0}+p_{2S+1})}+\frac{p_{4S+1}(a_{2S}p_{4S+1}+p_{2S}p_{4S})}{(p_{2S}+p_{4S+1})(p_{4S}+p_{4S+1})}+\frac{1}{2}\sum_{k=1}^{2S-1}\Bigg{[}\frac{(a_{k}p_{k+1}+p_{k}p_{2S+k})p_{k+1}}{(p_{k}+p_{k+1})(p_{k+1}+p_{2S+k})}$ (25)
@@
$+\frac{(a_{k}p_{2S+1+k}+p_{k}p_{2S+k})p_{2S+1+k}}{(p_{k}+p_{2S+1+k})(p_{2S+k}+p_{2S+1+k})}\Bigg{]}\Bigg{\}}.$
Remarkably, both measures ${\cal U}$ and ${\cal F}$ for axially symmetric states involve only square radicals and are expressed directly in terms of the matrix elements of the density matrix (2) and its eigenvalues (3). Draw attention that the $2S$ off-diagonal entries $u_{k}$ appear everywhere under the modulus symbol.
In contrast to Eqs. (8), (9), (19) and (20), algebraic expressions (13), (14), (24) and (25) for the branches of quantum correlations are surprisingly simple and very convenient for programming on a computer.
IV System at thermal equilibrium
As an application of the derived formulas, we consider spin-(1/2,$S$) system in the state of thermal equilibrium. Take the Hamiltonian $\cal H$ commuting with ${\cal S}_{z}$. It will have an axially symmetrical structure (portrait) like (2) and can be written as